1. At what temperature will a solution that is composed of 0.73 moles of glucose (C6H12O6) in 650mL of water (H2O) begin to boil?
1
Expert's answer
2021-10-25T03:50:15-0400

Using the relation ΔTb​=mKb


Molarity "=\frac{0.73×1000}{650}=1.123M"


Hence, "\u0394Tb=1.123\u00d70.52=0.58K"


Hence, the boiling point of the solution will increase by 0.58 K


New boiling point "=373.15+0.58=373.73 k"